OpenStudy (anonymous):

Angles 1 and 2 are complementary. IF the measure of angle 1 is 3 degrees less than twice the measure of angle 2, what is the measure of angle 1?

OpenStudy (anonymous):

angle 1+angle 2 = 90 lets say angle 2 = x then angle 1 would = 2x-3 (three less than twice x) so plug it in the equation 2x-3+x = 90 simplify to 3x-3 = 90 add 3 to both sides 3x = 93 divide both sides by 3 x = 31 so we said angle 2 is x, so plug x in to angle 1's formula 2(31)-3 = angle 1 62-3 = angle 1 angle 1 = 59 degrees TAH DAH! MATH!

OpenStudy (anonymous):

doesn't look like u hate algebra so much... :)

OpenStudy (anonymous):

THANK YOU SO MUCH. I GOT IT!